My book says: write an equation in standard form Ax + By=C of the line that satisfies the given conditions. Use integer values for A,b,c.
Then the problem is: m=4/5, passes through(2,3)
Here's what i did but i didn't and can't solve it to find the right answer.
y-3=4/5(x-2)
y-3=4/5x-8/5
that's all i really got

Found 2 solutions by solver91311, jim_thompson5910:
Answer by solver91311(24713)



You are right where you want to be.

Multiply both sides of your equation by 5



Add to both sides



Add to both sides



Done.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by jim_thompson5910(35256)

y-3 = 4/5x-8/5

5y-15 = 4x-8 ... Multiply EVERY term by LCD by 5.

5y = 4x-8+15

5y = 4x+7

5y-4x = 7

-4x+5y = 7

-1(-4x+5y) = -1*7

4x - 5y = -7

So the answer is 4x - 5y = -7